OpenStudy (anonymous):

What is parallel to 3x-5y=7 and passes through (0,6)?

OpenStudy (anonymous):

U WANT EQUATION?

OpenStudy (zzr0ck3r):

ax+by = c has slope -a/b and we know that something is parallel if it has the same slope.

OpenStudy (anonymous):

\[3x-5y=7\]\[-5y=7-3x\]\[y=\frac{3}{5}x-\frac{7}{5}\] A parallel function has the same slope, \[y=\frac{3}{5}x-b\] Since it goes through (0,6) you plug in the points to find b,\[(6)=\frac{3}{5}(0)-b\]\[6=-b\]\[b=-6\] Therefore, the function that is parallel to the original one and passes through (0,6) is,\[y=\frac{3}{5}x-6\]

OpenStudy (anonymous):

THANK YOU GUYS!!!!!!!